OpenStudy (anonymous):

Your fixed expenses are $1,763.25/month. You saved 6 months' worth for an emergency fund in a savings account earning a 4.5% APR over 3 years. After 3 years, you withdrew $4,360.00 because of losing your job. What is your balance after the withdrawal?

OpenStudy (anonymous):

$7,712.97 <--That is the answer I got. Just seeing if anyone could check it.

jimthompson5910 (jim_thompson5910):

You have 1763.25*6 = 10,579.50 dollars to invest

jimthompson5910 (jim_thompson5910):

You invest it for 3 years at 4.5% (simple I'm assuming) interest so you gain 10579.50*(0.045)*3 = 1428.23 dollars in interest alone over 3 years

jimthompson5910 (jim_thompson5910):

which means you have 10579.50+1428.23 = 12007.73 dollars in the account total after 3 years

jimthompson5910 (jim_thompson5910):

Now withdraw 4360 to get 12007.73 - 4360 = 7647.73 So you were close